OpenStudy (anonymous):

How fast (metric units) is a ball going after 3.0 seconds if you drop it off a very high building? It would be 9.8 m/s^2 since gravity is a constant speed, right?

OpenStudy (anonymous):

If there is no external force acting on a free falling body then it must fall with the speed of 9.8 m/s^2(acceleration due to gravity) irrespective of its mass or any physical quantity.

OpenStudy (anonymous):

Acceleration due to gravity is always different on different planets. which depends on the density of the celestial body.

OpenStudy (anonymous):

So the answer wouldn't be 29.4 m/sec^2?

OpenStudy (anonymous):

or actually -29.4 m/sec^2

OpenStudy (anonymous):

v=v0-gt v0 = 0 v = -gt v = -9.8 * 3 = -29.4 m/s ???

OpenStudy (anonymous):

Since they just want how fast and not which direction, you can ignore the - sign, so just 29.4m/s

OpenStudy (anonymous):

for a free fall \[u=0\] \[g=9.8 m/s ^{2}\] putting this value in v=u+gt we get \[v=0+9.8*3 \] (since time is 3 second) therefore v=29.4\[m/s\]